# Spectral Hardening Reveals Afterglow Emergence in Long-Duration Fast X-ray Transients: A Case Study of GRB 250404A/EP250404a

### v1

@@ Abstract
The prompt emission and afterglow phases of gamma-ray bursts (GRBs) have been extensively studied, yet the transition between these two phases remains inadequately characterized due to limited multiwavelength observational coverage. Among the recent growing samples of fast X-ray transients observed by Einstein Probe (EP), a subgroup of gamma-ray bursts are captured with long-duration X-ray emission, potentially containing featured evolution from prompt emission to the afterglow phase. In this Letter, we present a detailed analysis of GRB 250404A/EP250404a, a bright fast X-ray transient detected simultaneously by EP and Fermi/GBM in X-rays and gamma-rays. Its continuous X-ray emission reveals a long-duration tail, accompanied by distinct spectral evolution manifested by the spectral index $\alpha_{\rm X}$ with an initial softening, followed by an evident hardening, eventually reaching a plateau at the value of $\sim$ -2. Early optical and near-infrared observations enable broadband modeling with forward- and reverse-shock components, confirming that the X-ray hardening signals the emergence of the external-shock afterglow. From this spectral hardening we infer that the prompt phase in soft X-rays lasted $\sim 300\;\mathrm{s}$, which is more than three times longer than the gamma-ray $T_{90}$. This well-tracked soft-hard-flat spectral pattern provides a clear indication of afterglow emergence from the fading prompt emission and offers a practical criterion for identifying a distinct population of GRBs among fast X-ray transients, even when the detection of the gamma-ray counterpart or obvious temporal break is absent.
Transient sources; High energy astrophysics; X-ray transient sources; Gamma-ray bursts
@@ 1 Introduction
The afterglow phase of a gamma-ray burst (GRB) is conventionally defined as the stage that follows the end of the prompt emission (Zhang, 2018). From a theoretical perspective, the prompt emission is attributed to internal energy dissipation within the relativistic jet (Rees & Meszaros, 1994; Paczynski & Xu, 1994; Sari & Piran, 1997; Zhang et al., 2006), while the afterglow arises from the interaction between the ejecta and the circumburst medium (Paczynski & Rhoads, 1993; Mészáros & Rees, 1997). Observationally, the prompt emission typically exhibits rapid emission with erratic variability in the sub-MeV range, which is detected by GRB-triggering instruments (Zhang, 2014). In contrast, the afterglow is distinguished by broad-band spectra following a broken power-law shape and light curves displaying multi-segment broken power-law behavior, usually captured through multiwavelength follow-up observations (Sari et al., 1998). Extensive studies have explored the rich observational features of prompt emission, including its duration (Kouveliotou et al., 1993), spectral components (Zhang et al., 2011), and spectral evolution (Medvedev, 2006; Gao et al., 2021). Meanwhile, significant efforts have been devoted to understanding the afterglow phase, which is generally well-explained by synchrotron radiation originating from the external forward shock (FS) or the external reverse shock (RS) during the reverse shock crossing stage (Sari et al., 1998; Zhang & Kobayashi, 2005; Zhang et al., 2003).
@@
However, the transition from the prompt emission phase to the afterglow phase is often not fully traced owing to a combination of instrumental, observational, and physical challenges. Instrumentally, the narrow field of view of most telescopes, except for GRB-triggering detectors, limits opportunities for obtaining simultaneous multiwavelength coverage during the prompt emission. Follow-up observations intended to capture the afterglow often begin with delays, resulting in incomplete coverage of the early emission period. Observationally, the curvature effect at the end of the prompt phase causes the emission to progressively soften and fade over time (Fenimore et al., 1996; Kumar & Panaitescu, 2000; Dermer, 2004), shifting it outside the energy range or below the sensitivity threshold of gamma-ray detectors. Physically, in the initial hours after the burst, the observed emission may be a complex superposition of multiple components (Zhang et al., 2006), including internal dissipation from late central engine activity, emission from the external RS, and from the external FS. As a result, it remains challenging to pinpoint when the afterglow emission from external shocks starts to appear and eventually dominates following the decay of the prompt emission.
@@
The Einstein Probe (EP) Mission (Yuan et al., 2025), with its wide field of view, has detected dozens of fast X-ray transients, several of which exhibit prompt emission counterparts in gamma-rays (Yin et al., 2024; Liu et al., 2025; Jiang et al., 2025b), confirming their origin from the GRB internal energy dissipation of the relativistic jet. Notably, most of the EP-detected GRBs display soft X-ray tails that extend well beyond the duration of the gamma-ray emission, offering a valuable opportunity to study the spectral evolution during the late prompt emission and the transition to the afterglow emission. Crucially, the onboard trigger of EP/Wide-field X-ray Telescope (WXT) enables rapid localization with an accuracy of a few arcminutes, which is further refined to several arcseconds by the automatic follow-up observation from EP/Follow-up X-ray Telescope (FXT) within minutes. This capability facilitates timely multiwavelength follow-up observations by ground-based observatories during the early afterglow phase, concurrent with ongoing X-ray detection. Such synergy yields critical broadband insights into the distinct emission components. As a representative example, in this Letter, we present the EP-triggered detection of GRB 250404A/EP250404a featuring rich observations coverage X-ray and gamma-ray prompt emission, as well as optical and near-infrared follow-ups within few hours. A fast flux rise is observed in both X-ray and optical bands, following the spectral hardening in the X-ray emission. We performed the afterglow fitting on the multiwavelength data with a combined FS and RS model, and identified the very early afterglow emergence marked by clear spectral evolution in the X-ray band.
@@ 2.1 Einstein Probe Observations
The WXT cleaned event files and response files were generated following the standard data reduction pipelines implemented in the WXT Data Analysis Software (WXTDAS v2.10; Y. Liu et al., in preparation) and the calibration database (CALDB, v1.0; H.-Q. Cheng et al., in preparation). Source photons were extracted from a circular region with a radius of 999’9 ’, while background photons were extracted from an annular region with inner and outer radii of 181818’18 ’ and 363636’36 ’, respectively.
@@
We note that WXT underwent a significant slew (indicated by the dashed curve in the first panel on the left of Figure 1) during the main emission phase. As a result, we caution that the WXT data presented in this work should be treated as qualitative references rather than precise measurements. The WXT light curve shown in Figure 1 is therefore intended to illustrate the temporal coverage provided by EP. Our quantitative analysis of the burst properties relies primarily on the FXT data.
@@ 2.2 Fermi/GBM Observations
The Fermi Gamma-ray Burst Monitor (GBM; Meegan et al., 2009) detected GRB 250404A (Mukherjee et al., 2025) at $T_{0}$, with a calculated location consistent with that of EP250404a (Fermi GBM Team, 2025). We retrieved the time-tagged event (TTE) data set covering the time range of GRB 250404A using the Python package heapy [^1] https://github.com/jyangch/heapy. From the 12 sodium iodide detectors onboard, we selected detectors n0, n1, and n9, which had the smallest viewing angles relative to the GRB source direction. In addition, we included the brightest bismuth germanium oxide detector, b0, to extend the energy coverage. Data reduction and analysis were performed with heapy, following the standard procedures described in Zhang et al. (2011) and Yang et al. (2022).
@@ 2.3 Ground-based Observations
We conducted ground-based follow-up observations in the optical and near-infrared bands using the Multi-channel Photometric Survey Telescope (Mephisto), the Tsinghua-Nanshan Optical Telescope (TNOT), the Sun Yat-sen University 80 cm telescope (SYSU 80cm), the Burst Observer and Optical Transient Exploring System (BOOTES)-4/María Eva Telescope (MET), and the Schmidt telescope. Additional imaging data were obtained from the the Alhambra Faint Object Spectrograph and Camera (ALFOSC) mounted on the 2.56 m Nordic Optical Telescope (NOT), located at the Roque de los Muchachos Observatory, La Palma, Spain, the Half-Meter telescope (HMT) located at Nanshan Observatory, Xinjiang, China, the JinShan 0.5 m (50D) and 1 m (100C) telescopes located at Altay Observatory, Xinjiang, China. Spectroscopic observations were carried out with the 2.4-meter telescope at the Yunnan Observatories (GMG-2.4m), covering the wavelength range of 3600–7460 Å. A summary of these observations is provided in Figure 3, Table LABEL:tab:photoag, Table LABEL:tab:photoclear and Figure 6.
@@ 2.3.1 Photometry
The optical counterpart of GRB 250404A/EP250404a was first detected by HMT at 14:24:59 UT on 2025 April 4, 313 s after the $T_{0}$. It was located at (J2000) R.A. = $08^{\rm hr}20^{\rm m}14.54^{\rm s}$, decl. = $+35^{\circ}31^{\prime}41.57^{\prime\prime}$ with an uncertainty of $0.5^{\prime\prime}$ (Jiang et al., 2025a). The observation lasted approximately 2.6 hours and consisted of a series of unfiltered exposures. We calibrated the photometry of the HMT data with Gaia DR3 $G$-band reference stars.
@@
At $\sim$ 734 s after the $T_{0}$, Mephisto (Yang et al., 2024; Chen et al., 2024; Du et al., 2025) started the observation in the $u$, $v$, $g$, $r$, $i$ and $z$ bands. The optical afterglow was clearly detected at R.A. = 125.0607, decl. = 35.5282 (J2000). The $u$, $g$ and $i$ bands exhibited a rapid flux rise within the first $\sim$ 300 s of the observation, followed by a decay phase characterized by differing slopes between the early and late stages. Notably, the measurements in the blue bands of Mephisto deviate from the predictions of the standard afterglow model considering the correction for Galactic extinction along the line of sight (see Figure 4). This indicates the presence of additional extinction effects from the host galaxy or the surrounding medium.
@@
We obtained a sequence of images in the $g,\,r,\,i,$ and $z$ bands using the 50D and 100C telescopes located at Altay Observatory. These observations were carried out between 0.8 and 4.6 hours after the $T_{0}$. Additionally, NOT/ALFOSC conducted follow-up observations in the same four filters approximately 8.3 hours after the $T_{0}$.
@@
At $\sim$ $T_{0}$ + 1214 s, we also observed the field of the GRB 250404A/EP250404a with the 0.80m TNOT at Nanshan Station of Xinjiang Astronomical Observatory. A series of $g$, $r$ and $i$ band images were obtained, clearly detecting the optical afterglow at R.A. = 125.0606, decl. = 35.5282 (J2000). The fluxes derived from TNOT observations are in agreement with those measured by Mephisto in the corresponding optical bands.
@@
In the $J$ band, Sun Yat-sen University 80cm infrared telescope started the observation at $\sim$ $T_{0}$ + 4021 s, obtaining 183 exposures of 20 s each. A counterpart was detected at the position of the optical afterglow in the stacked images.
@@
Additional observations using clear filters have also captured the long-term decay features of the afterglow. Following the EP trigger, the 0.6m BOOTES-4/MET robotic telescope at Lijiang Astronomical Observatory automatically responded to this event at $\sim$ $T_{0}$ + 1203 s. A series of clear-filter images were collected, revealing a source consistent with the optical afterglow position. The magnitudes were measured using Gaia DR3 $G$-band as the reference. Furthermore, the 0.9 m Schmidt telescope at Xinglong Observatory, monitored the source over an extended period from $T_{0}$ + 1251 s to $T_{0}$ + 6379 s, capturing a series of unfiltered images. The magnitudes for these observations were calibrated using Gaia DR2 $G$-band as the reference.
@@ 2.3.2 Spectroscopy
The spectroscopic observation was conducted using the GMG-2.4m telescope with Grism #14, which provides a wavelength coverage of 3600-7460 Å(Wang et al., 2019; Xin et al., 2020). A single spectrum with an exposure time of 1800 s was obtained, beginning at $T_{0}$ + 3519 s. The spectrum was reduced using standard IRAF procedures (v2.18.1; Tody, 1986, 1993; Observatories & community, 2025). We identified multiple metal absorption features as shown in Figure 6, including Si IV at 1394 Å, Si IV at 1403 Å, Si II at 1527 Å, C IV at 1549 Å, Fe II at 1608 Å, Al II at 1671 Å, Al III at 1855 Å, Al III at 1863 Å, Zn II/Cr II at 2026 Å, Zn II/Cr II at 2062 Å, Fe II at 2344 Å, Fe II at 2374 Å and Fe II at 2383 Å. These features consistently indicate a redshift of z $\sim$ 1.88.

### v2

@@ Abstract
The prompt emission and afterglow phases of gamma-ray bursts (GRBs) have been extensively studied, yet the transition between these two phases remains inadequately characterized due to limited multiwavelength observational coverage. Among the recent growing samples of fast X-ray transients observed by Einstein Probe (EP), a subgroup of GRBs are captured with long-duration X-ray emission, potentially containing featured evolution from prompt emission to the afterglow phase. In this Letter, we present a detailed analysis of GRB 250404A/EP250404a, a bright fast X-ray transient detected simultaneously by EP and the Fermi Gamma-ray Burst Monitor in X-rays and gamma rays. Its continuous X-ray emission reveals a long-duration tail, accompanied by distinct spectral evolution manifested by the spectral index $\alpha_{\rm X}$ with an initial softening, followed by an evident hardening, eventually reaching a plateau at the value of $\sim$ -2. Early optical and near-infrared observations enable broadband modeling with forward- and reverse-shock components, confirming that the X-ray hardening signals the emergence of the external-shock afterglow. From this spectral hardening we infer that the prompt phase in soft X-rays lasted $\sim 300\;\mathrm{s}$, which is more than 3 times longer than the gamma-ray $T_{90}$. This well-tracked soft-hard-flat spectral pattern provides a clear indication of afterglow emergence from the fading prompt emission and offers a practical criterion for identifying a distinct population of GRBs among fast X-ray transients, even when the detection of the gamma-ray counterpart or obvious temporal break is absent.
@@ 1 Introduction
The afterglow phase of a gamma-ray burst (GRB) is conventionally defined as the stage that follows the end of the prompt emission (B. Zhang, 2018). From a theoretical perspective, the prompt emission is attributed to internal energy dissipation within the relativistic jet (M. J. Rees & P. Meszaros, 1994; B. Paczynski & G. Xu, 1994; R. Sari & T. Piran, 1997; B. Zhang et al., 2006), while the afterglow arises from the interaction between the ejecta and the circumburst medium (B. Paczynski & J. E. Rhoads, 1993; P. Mészáros & M. J. Rees, 1997). Observationally, the prompt emission typically exhibits rapid emission with erratic variability in the sub-MeV range, which is detected by GRB-triggering instruments (B. Zhang, 2014). In contrast, the afterglow is distinguished by broadband spectra following a broken power-law shape and light curves displaying multisegment broken power-law behavior, usually captured through multiwavelength follow-up observations (R. Sari et al., 1998). Extensive studies have explored the rich observational features of prompt emission, including its duration (C. Kouveliotou et al., 1993), spectral components (B.-B. Zhang et al., 2011), and spectral evolution (M. V. Medvedev, 2006; H.-X. Gao et al., 2021). Meanwhile, significant efforts have been devoted to understanding the afterglow phase, which is generally well explained by synchrotron radiation originating from the external forward shock (FS) or the external reverse shock (RS) during the RS crossing stage (R. Sari et al., 1998; B. Zhang & S. Kobayashi, 2005; B. Zhang et al., 2003).
@@
However, the transition from the prompt emission phase to the afterglow phase is often not fully traced owing to a combination of instrumental, observational, and physical challenges. Instrumentally, the narrow field of view of most telescopes, except for GRB-triggering detectors, limits opportunities for obtaining simultaneous multiwavelength coverage during the prompt emission. Follow-up observations intended to capture the afterglow often begin with delays, resulting in incomplete coverage of the early emission period. Observationally, the curvature effect at the end of the prompt phase causes the emission to progressively soften and fade over time (E. E. Fenimore et al., 1996; P. Kumar & A. Panaitescu, 2000; C. D. Dermer, 2004), shifting it outside the energy range or below the sensitivity threshold of gamma-ray detectors. Physically, in the initial hours after the burst, the observed emission may be a complex superposition of multiple components (B. Zhang et al., 2006), including internal dissipation from late central engine activity and emission from the external RS and the external FS. As a result, it remains challenging to pinpoint when the afterglow emission from external shocks starts to appear and eventually dominates following the decay of the prompt emission.
@@
The Einstein Probe (EP) mission (W. Yuan et al., 2025), with its wide field of view, has detected dozens of fast X-ray transients, several of which exhibit prompt emission counterparts in gamma rays (Y.-H. I. Yin et al., 2024; Y. Liu et al., 2025; S.-Q. Jiang et al., 2025), confirming their origin from the GRB internal energy dissipation of the relativistic jet. Notably, most of the EP-detected GRBs display soft X-ray tails that extend well beyond the duration of the gamma-ray emission, offering a valuable opportunity to study the spectral evolution during the late prompt emission and the transition to the afterglow emission. Crucially, the onboard trigger of the EP Wide-field X-ray Telescope (WXT) enables rapid localization with an accuracy of a few arcminutes, which is further refined to several arcseconds by the automatic follow-up observation from the EP Follow-up X-ray Telescope (FXT) within minutes. This capability facilitates timely multiwavelength follow-up observations by ground-based observatories during the early afterglow phase, concurrent with ongoing X-ray detection. Such synergy yields critical broadband insights into the distinct emission components. As a representative example, in this Letter, we present the EP-triggered detection of GRB 250404A/EP250404a featuring rich observations covering X-ray and gamma-ray prompt emission, as well as optical and near-infrared follow-ups within a few hours. A fast flux rise is observed in both X-ray and optical bands, following the spectral hardening in the X-ray emission. We performed the afterglow fitting on the multiwavelength data with a combined FS and RS model and identified the very early afterglow emergence marked by clear spectral evolution in the X-ray band.
@@ 2.1 EP Observations
The WXT cleaned event files and response files were generated following the standard data reduction pipelines implemented in the WXT Data Analysis Software (WXTDAS v2.10; Y. Liu et al., in preparation) and the calibration database (CALDB, v1.0; H. Cheng et al., 2025). Source photons were extracted from a circular region with a radius of 99’, while background photons were extracted from an annular region with inner and outer radii of $18$’ and $36$’, respectively.
@@
We note that WXT underwent a significant slew (indicated by the dashed curve in the first panel on the left of Figure 1) during the main emission phase. As a result, we caution that the WXT data presented in this work should be treated as qualitative references rather than precise measurements. The WXT light curve shown in Figure 1 is therefore intended to illustrate the temporal coverage provided by EP. Our quantitative analysis of the burst properties (see Tables 1, 3 and 4) relies primarily on the FXT data.
@@ 2.2 Fermi/GBM Observations
The Fermi Gamma-ray Burst Monitor (GBM; C. Meegan et al., 2009) detected GRB 250404A (O. Mukherjee et al., 2025) at $T_{0}$, with a calculated location consistent with that of EP250404a (Fermi GBM Team, 2025). We retrieved the time-tagged event data set covering the time range of GRB 250404A using the Python package heapy [^1] https://github.com/jyangch/heapy. From the 12 sodium iodide detectors on board, we selected detectors n0, n1, and n9, which had the smallest viewing angles relative to the GRB source direction. In addition, we included the brightest bismuth germanium oxide detector, b0, to extend the energy coverage. Data reduction and analysis were performed with heapy, following the standard procedures described in B.-B. Zhang et al. (2011) and J. Yang et al. (2022).
@@ 2.3 Ground-based Observations
We conducted ground-based follow-up observations in the optical and near-infrared bands using the Multi-channel Photometric Survey Telescope (Mephisto), the Tsinghua-Nanshan Optical Telescope (TNOT), the Sun Yat-sen University 80 cm telescope (SYSU 80cm), the Burst Observer and Optical Transient Exploring System (BOOTES)-4/María Eva Telescope (MET), and the 60/90 cm Schmidt telescope. Additional imaging data were obtained from the the Alhambra Faint Object Spectrograph and Camera (ALFOSC) mounted on the 2.56 m Nordic Optical Telescope (NOT), located at the Roque de los Muchachos Observatory, La Palma, Spain; the Half-Meter telescope (HMT) located at Nanshan Observatory, Xinjiang, China; and the JinShan 0.5 m (50D) and 1 m (100C) telescopes located at Altay Observatory, Xinjiang, China. Spectroscopic observations were carried out with the 2.4 m telescope at the Yunnan Observatories (GMG-2.4 m), covering the wavelength range of 3600–7460 Å. A summary of these observations is provided in Figures 2 and LABEL:tab:photoag and Tables LABEL:tab:photoclear and Figure 6.
Figure 2: Multiwavelength observations of GRB 250404A/EP250404a and afterglow modeling with the best-fit parameters from the FS+RS model with $u$-, $v$-, and $g$-band correction factors. The multiwavelength data utilized in the afterglow fitting are marked with filled points, while the optical observations on clear filters that are not included in the afterglow fitting are marked with open points. The optical and near-infrared data have been corrected for Galactic extinction, which is E(BV)=0.073E(B-V)=0.073 mag (E. F. Schlafly & D. P. Finkbeiner, 2011). We assume a total to selective extinction ratio of $R_{V}=3.1$ according to the extinction law from E. L. Fitzpatrick (1999). The inverted triangle points signify upper limits. The best-fit FS+RS, FS, and RS models are shown with solid, dashed-dotted and dotted lines, respectively. The dark blue dashed vertical lines correspond to the same time interval where the spectra start to harden and reach a plateau.
@@ 2.3.1 Photometry
The optical counterpart of GRB 250404A/EP250404a was first detected by HMT at 14:24:59 UT on 2025 April 4, 313 s after the $T_{0}$. It was located at (J2000) R.A. = $08^{\rm hr}20^{\rm m}14.54^{\rm s}$, decl. = $35^{\circ}31^{\prime}41.57^{\prime\prime}$ with an uncertainty of $0.5^{\prime\prime}$ (S. Q. Jiang et al., 2025). The observation lasted approximately 2.6 hr and consisted of a series of unfiltered exposures. We calibrated the photometry of the HMT data with Gaia DR3 $G$-band reference stars.
@@
At $\sim$ 734 s after the $T_{0}$, Mephisto (Y.-P. Yang et al., 2024; X. Chen et al., 2024; G. Du et al., 2025) started the observation in the $u$, $v$, $g$, $r$, $i$ and $z$ bands. The optical afterglow was clearly detected at R.A. = 125.0607, decl. = 35.5282 (J2000). The $u$, $g$ and $i$ bands exhibited a rapid flux rise within the first $\sim$ 300 s of the observation, followed by a decay phase characterized by differing slopes between the early and late stages. Notably, the measurements in the blue bands of Mephisto deviate from the predictions of the standard afterglow model considering the correction for Galactic extinction along the line of sight (see Figure 3). This indicates the presence of additional extinction effects from the host galaxy or the surrounding medium.
@@
We obtained a sequence of images in the g,r,i,g,\,r,\,i, and $z$ bands using the 50D and 100C telescopes located at Altay Observatory. These observations were carried out between 0.8 and 4.6 hr after the $T_{0}$. Additionally, NOT/ALFOSC conducted follow-up observations in the same four filters approximately 8.3 hr after the $T_{0}$.
@@
At $\sim$ $T_{0}$ + 1214 s, we also observed the field of the GRB 250404A/EP250404a with the 0.80m TNOT at Nanshan Station of Xinjiang Astronomical Observatory. A series of $g$-, $r$- and $i$-band images were obtained, clearly detecting the optical afterglow at R.A. = 125.0606, decl. = 35.5282 (J2000). The fluxes derived from TNOT observations are in agreement with those measured by Mephisto in the corresponding optical bands.
@@
In the $J$ band, the Sun Yat-sen University 80cm infrared telescope started the observation at $\sim$ $T_{0}$ + 4021 s, obtaining 183 exposures of 20 s each. A counterpart was detected at the position of the optical afterglow in the stacked images.
@@
Additional observations using clear filters have also captured the long-term decay features of the afterglow. Following the EP trigger, the 0.6m BOOTES-4/MET robotic telescope at Lijiang Astronomical Observatory automatically responded to this event at $\sim$ $T_{0}$ + 1203 s. A series of clear-filter images were collected, revealing a source consistent with the optical afterglow position. The magnitudes were measured using Gaia DR3 $G$ band as the reference. Furthermore, the 60/90 cm Schmidt telescope at Xinglong Observatory monitored the source over an extended period from $T_{0}$ + 1251 s to $T_{0}$ + 6379 s, capturing a series of unfiltered images. The magnitudes for these observations were calibrated using Gaia DR2 $G$ band as the reference.
@@ 2.3.2 Spectroscopy
The spectroscopic observation was conducted using the GMG-2.4m telescope with grism #14, which provides a wavelength coverage of 3600-7460 Å(C.-J. Wang et al., 2019; Y.-X. Xin et al., 2020). A single spectrum with an exposure time of 1800 s was obtained, beginning at $T_{0}$ + 3519 s. The spectrum was reduced using standard IRAF procedures (v2.18.1; D. Tody, 1986, 1993; N. O. A. Observatories & T. I. community, 2025). We identified multiple metal absorption features as shown in Figure 6, including Si IV at 1394 Å, Si IV at 1403 Å, Si II at 1527 Å, C IV at 1549 Å, Fe II at 1608 Å, Al II at 1671 Å, Al III at 1855 Å, Al III at 1863 Å, Zn II/Cr II at 2026 Å, Zn II/Cr II at 2062 Å, Fe II at 2344 Å, Fe II at 2374 Å and Fe II at 2383 Å. These features consistently indicate a redshift of $z$ $\sim$ 1.88.
Figure 3: X-ray/optical/near-infrared afterglow specific flux density spectra of GRB 250404A/EP250404a in different time intervals. The best-fit afterglow model is indicated by dashed-dotted gray lines. The red curves correspond to the partial host galaxy extinction curve derived from the best-fit correction factors from the afterglow modeling. For comparison, the yellow, blue, and green curves show the average extinction laws of the MW, SMC, and LMC, respectively, assuming a reddening value of E(BV)=0.1E(B-V)=0.1 mag.